OpenStudy (anonymous):

Find the derivative of the function. y = (e^(7u) − e^(−7u)/(e^(7u) + e^(−7u))

OpenStudy (anonymous):

how did you get that?

OpenStudy (loser66):

don't you see the form of sinh (7u)/cosh(7u)??

OpenStudy (anonymous):

I thought that the quotient rule would be used.

OpenStudy (loser66):

you have the formula of \(sinh x = \dfrac{e^x -e^{-x}}{2}\) and \(cosh x = \dfrac{e^x+e^{-x}}{2}\) yours is 7u, not x . and combine them, you have sinh (7u)/ cosh (7u) = tanh (7u) derivative of tanh(7u)= sech (7u) that's it

OpenStudy (loser66):

nope?? don't agree with me?? hehehe.. if you want to apply quotient rule, ok, go ahead.